# ]https://zf-wei.github.io

### v2

@@ III.4 Results on real-world networks
Figure 16: Mean ECS for various methods over the percentage of selected nodes on the two real-world networks with $16$-dimensional embeddings. (a) and (b) use random node selection while (c) and (d) use targeted node selection. (a) and (c) corresponds to the email-EU-core network while (b) and (d) corresponds to the AS network.
IV Conclusion
@@
Our experimental findings revealed a general decline in community similarity as more nodes were selected and their adjacent edges deleted. Notably, targeted node selection led to a more pronounced and rapid decline in community similarity compared to random node selection. Analysis of clustering similarity scores suggested that LFR networks with lower mixing parameters, indicative of a stronger community structure, exhibit greater robustness to network perturbation. Moreover, diverse graph embedding methods displayed varying degrees of robustness in network community detection. Specifically, LLE and LE demonstrated proficiency within the family of matrix-based methods. For smaller real-world networks with a limited number of vertices, the more robust embedding method is LLE with a modest embedding dimension, typically on the order of tens — such as the 16 dimensions employed in our experiment. In the case of larger real-world networks, the LE embedding method showed greater robustness to partition. Specifically, a higher embedding dimension is preferable when the community structure is less pronounced, whereas a lower embedding dimension is more effective when the network exhibits a stronger community structure. Within the family of random walk methods, our results consistently highlighted node2vec as the superior choice. Notably, node2vec also outperformed all methods within the family of matrix factorization methods.
@@
In contemplating potential avenues for future research, it is noteworthy that our current experiments did not involve parameter optimization in the graph embedding algorithms, primarily due to its inherent computational expense. To attain a more exhaustive and comprehensive insight into the robustness of community detection using graph embedding methods, we may explore the incorporation of parameter optimization in our forthcoming studies.
Appendix A Experimental results using $32$-dimensional embedding
In this appendix, we present figures on experimental results using embedding dimension $32$. In Figs. 17, 18, 19, 20, 21, 22 and 23, subfigures in the top row, i.e., (a) and (b), correspond to random node selection. More specifically, subfigure (a) corresponds to experimental results on networks with $1,000$ nodes and subfigure (b) depicts experimental results on networks with $10,000$ nodes. Similarly, subfigures in the bottom row, i.e., (c) and (d), correspond to targeted node selection. More specifically, subfigure (c) corresponds to experimental results on networks with $1,000$ nodes and subfigure (d) depicts experimental results on networks with $10,000$ nodes.
We notice that some embedding methods exhibit different robustness with varying embedding dimensions. For instance, when using the Laplacian eigenmap method, in networks with strong community partitions (i.e., $\mu=0.01$ or $\mu=0.1$), the embedding dimension significantly impacts the decay rate of similarity, regardless of the node selection strategy. In Fig. 2(a), with an embedding dimension of $16$ and random node selection, the curve for $\mu=0.1$ remains stable around an ECS value of approximately $1.0$ until $60\%$ of nodes are selected. Between $60\%$ and $85\%$ of nodes selected, the curve gradually decreases from $1.0$ to $0.84$ (a $16\%$ decrease). In contrast, in Fig. 17(a), with a $32$-dimensional embedding, the curve for $\mu=0.1$ remains stable around $1.0$ only until $25\%$ of nodes are selected. Then, between $25\%$ and $85\%$ of nodes selected, the curve sharply decreases from $1.0$ to $0.53$ (a $47\%$ decrease). For the HOPE embedding method, our results suggest that a $32$-dimensional embedding (i.e., Fig. 19) yields higher ECS scores than a $16$-dimensional embedding (i.e., Fig. 6). Specifically, the data points on the green curve (corresponding to $\mu=0.4$) in Fig. 19(a) have higher ECS scores than those on the green curve in Fig. 6(a) at respective percentages of selected nodes.
Figure 17: Mean ECS using LE method over the percentage of selected nodes on LFR benchmark graphs with $32$-dimensional embeddings. Random node selection is used in (a) and (b), while targeted selection is used in (c) and (d). (a) and (c) are for networks with $1,000$ nodes, and (b) and (d) are for networks with $10,000$ nodes.
Figure 18: Mean ECS using LLE method over the percentage of selected nodes on LFR benchmark graphs with $32$-dimensional embeddings. (a) and (b) apply random node selection, whereas (c) and (d) use targeted selection. Networks in (a) and (c) have $1,000$ nodes, and those in (b) and (d) have $10,000$ nodes.
Figure 19: Mean ECS using HOPE method over the percentage of selected nodes on LFR benchmark graphs with $32$-dimensional embeddings. While (a) and (b) use random node selection, (c) and (d) rely on targeted selection. Networks in (a) and (c) contain $1,000$ nodes, while those in (b) and (d) consist of $10,000$ nodes.
Figure 20: Mean ECS using M-NMF method over the percentage of selected nodes on LFR benchmark graphs with $32$-dimensional embeddings. (a) and (b) implement random node selection, while (c) and (d) apply targeted selection; (a) and (c) correspond to networks with $1,000$ nodes, and (b) and (d) to networks with $10,000$ nodes.
Figure 21: Mean ECS using DeepWalk method over the percentage of selected nodes on LFR benchmark graphs with $32$-dimensional embeddings. Random selection is employed in (a) and (b), while targeted selection is used in (c) and (d). (a) and (c) represent networks with $1,000$ nodes, and (b) and (d) represent those with $10,000$ nodes.
Figure 22: Mean ECS using LINE method over the percentage of selected nodes on LFR benchmark graphs with $32$-dimensional embeddings. (a) and (b) follow random node selection, whereas (c) and (d) adopt targeted node selection. Networks with $1,000$ nodes correspond to (a) and (c), and those with $10,000$ nodes correspond to (b) and (d).
Figure 23: Mean ECS using node2vec method over the percentage of selected nodes on LFR benchmark graphs with $32$-dimensional embeddings. Random node selection is applied in (a) and (b), while (c) and (d) use targeted selection. (a) and (c) deal with $1,000$-node networks, and (b) and (d) with $10,000$-node networks.
Figure 24: Mean ECS for various methods over the percentage of selected nodes on the two real-world networks with $32$-dimensional embeddings. (a) and (b) use random node selection while (c) and (d) use targeted node selection. (a) and (c) corresponds to the email-EU-core network while (b) and (d) corresponds to the AS network.
Appendix B LINE method with higher embedding dimensions
For the LINE embedding method, our analysis of networks with 1,000 nodes in the first column of Fig. 12 reveals a distinct trend: the curves exhibit a steep decline followed by a subsequent increase. According to our experiments, increasing the embedding dimensions in LINE is expected to mitigate the sharpness of this decline and increase, rendering the pattern less pronounced, as indicated in Fig. 25. However, for other embedding methods, a higher embedding dimension doesn’t necessarily lead to better performance with higher similarity scores. For instance, M-NMF with higher dimensions results in lower ECS scores, as shown in Fig. 26, compared to the first column of Fig. 8.
Figure 25: Mean ECS using LINE method over the percentage of selected nodes uniformly at random on LFR benchmark graphs with $1,000$ nodes. The first plot displays results using $128$-dimensional embeddings; the second plot displays results using $256$-dimensional embeddings.
Figure 26: Mean ECS using M-NMF method over the percentage of selected nodes uniformly at random on LFR benchmark graphs with $1,000$ nodes. The first plot displays results using $128$-dimensional embeddings; the second plot displays results using $256$-dimensional embeddings.

### v3

@@ IV.2.2 Large-Scale Information Network Embedding (LINE)
Figure 16: Mean ECS using node2vec method over the percentage of selected nodes on LFR benchmark graphs with 323232-dimensional embeddings. Random node selection is applied in (a) and (b), while (c) and (d) use targeted selection. (a) and (c) deal with 1,0001,0001 , 000-node networks, and (b) and (d) with 10,00010,00010 , 000-node networks.
In the case of node2vec, we consistently observe a decrease in ECS as we iteratively select additional nodes and remove their adjacent edges, which aligns with our intuitive expectations.
IV.2.4 Summary of Random Walk Methods
Similar to matrix factorization methods, the curve tends to exhibit higher values for smaller mixing parameters ($\mu$). Moreover, when using targeted node selection, the ECS shows a more pronounced and rapid decline compared to random node selection, as evident from the comparison between Fig. 15(a) and Fig. 15(c).
From our discussion of each method, we observe that DeepWalk may exhibit an unexpected increase while LINE experiences unexpected oscillations. In contrast, node2vec is free from these issues and consistently achieves the highest ECS similarity scores. See the Appendix A for a brief explanation on the advantage of Node2vec embedding in network community detection.
When considering the application of random walk embeddings for community detection, node2vec consistently proves to be a more robust choice over DeepWalk and LINE, across networks of varying sizes, the strength of the initial community partition, and perturbation type.
IV.3 Real-World Networks
In this section, we present the results of our experiments conducted on two real-world networks with labeled communities, i.e., the email-EU-core network and the AS network. The email-EU-core network is constructed using email data sourced from a prominent European research institution, comprising 986986986 nodes and 16,68716,68716 , 687 edges. For this network, we use the departmental affiliations of its members as the underlying community structure [77]. The AS network, derived from the AS Internet topology data collected in June 2009 by the Archipelago active measurement infrastructure, represents the inter-domain Internet topology with Autonomous Systems (ASes) as nodes and AS peerings as links, forming an AS-level topology graph. This dataset comprises 23,74823,74823 , 748 nodes and 58,41458,41458 , 414 edges. For this network, we use the densely connected AS groups as the underlying community structure [7]. We select them because they closely resemble power-law degree and community size distributions, as the synthetic networks counterpart (i.e., LFR benchmark graphs). To confirm the alignment of these real-world networks with the properties of LFR networks, we conducted statistical tests using the Python package powerlaw based on the statistical test derived in the work by Clauset et al [80]. The results indicate that both the email-EU-core network and the AS network exhibit power-law characteristics [81]. Specifically, in the email-EU-core network, the degree distribution adheres to a power-law distribution and the community size sequence follows an exponential distribution; while in the AS network, both degree and community size distributions conform to power laws.
[…]
We report experimental results using different methods with 161616-dimensional embedding on these two real-world networks in Fig. 17. The experimental results with 323232-dimensional embedding on the two real-world networks are depicted in Fig. 18.
Figure 17: Mean ECS for various methods over the percentage of selected nodes on the two real-world networks with 161616-dimensional embeddings. (a) and (b) use random node selection while (c) and (d) use targeted node selection. (a) and (c) corresponds to the email-EU-core network while (b) and (d) corresponds to the AS network.
Figure 18: Mean ECS for various methods over the percentage of selected nodes on the two real-world networks with 323232-dimensional embeddings. (a) and (b) use random node selection while (c) and (d) use targeted node selection. (a) and (c) corresponds to the email-EU-core network while (b) and (d) corresponds to the AS network.
[…]
As the number of removed selected nodes increases, the ECS sometimes rises, for example, in Fig. 17(a) and Fig. 18(a) for the M-NMF mehtod. To get more insights and mitigate this pathological increasing behavior, we have performed additional experiment and analyzed the ECS score of M-NMF with 128-dimensional embeddings. We notice that for clustering bigger networks where the expected number of communities is naturally higher, M-NMF required the use of a higher embedding dimension to effectively separate those communities. See the Appendix B for more details.
For the two real-world networks we analyzed, the “underlying community structure” is assigned based on the functional roles and real-world context of each network. While this assignment is somewhat artificial, we believe it is meaningful and provides a reasonable approximation. By leveraging these predefined community structures, we can still gain valuable insights into the robustness of our community detection methods using network embeddings.
IV.4 Quantitative Comparison of Graph Embedding Methods
[…]
$\Delta y=\frac{1}{13}\sum_{i=1}^{13}|y_{i}-y_{i-1}|.$
This quantity captures the average rate of decline in similarity as more nodes are selected and their adjacent edges removed. A smaller $\Delta y$ indicates that the embedding method better preserves the community structure under perturbation, i.e., it is more robust. Naturally, a more robust embedding method is characterized by having a larger $y_{0}$ (indicating better initial similarity) and a smaller $\Delta y$ (indicating better capacity to “maintain” community structure as more nodes are selected). Table 4 shows the numbers we got for the aforementioned parameters:
Table 4: Mean initial ECS values ($y_{0}$) and average decay rates ($\Delta y$) for different graph embedding methods under edge-deletion perturbations. Higher $y_{0}$ and lower $\Delta y$ indicate more robust community structure preservation. Results are based on networks with 10,00010{,}00010 , 000 nodes, embedding dimension 323232, and mixing parameter $\mu=0.1$.
Family Matrix Factorization Methods Random Walk Methods
Methods LE LLE HOPE M-NMF DeepWalk LINE node2vec
$y_{0}$ 0.914 1.000 0.566 0.943 0.853 0.881 1.000
$\Delta y$ 0.012 0.005 0.027 0.022 0.007 0.012 0.001
As observed from Table 4, for the parameter choices mentioned above, LLE and node2vec provide the most robust community detection results. This verifies our qualitative analysis. Additionally, we explore networks with varying node counts (1,0001,0001 , 000 or 10,00010,00010 , 000) and different mixing parameters ($\mu=0.01,0.1,0.2,0.3,0.4,0.5$), and we also conducted experiment with embedding dimensions of 161616 or 323232, for both random and targeted node selection. Given the variety of parameters, we compute for each embedding method the mean values over these configurations. We summarize the overall robustness of community detection across these seven graph embedding methods in Table 5.
Table 5: Mean robustness metrics for different embedding methods, averaged over multiple experimental settings including varying network sizes (1,0001{,}0001 , 000 and 10,00010{,}00010 , 000 nodes), mixing parameters ($\mu=0.01$ to 0.50.50.5), embedding dimensions (161616 and 323232), and both random and targeted node selection strategies.
Family Matrix Factorization Methods Random Walk Methods
Methods LE LLE HOPE M-NMF DeepWalk LINE node2vec
$y_{0}$ 0.896 0.912 0.651 0.766 0.837 0.877 0.921
$\Delta y$ 0.026 0.025 0.022 0.029 0.019 0.040 0.017
From these tables, we can confirm that LLE and node2vec provide the most robust community detection results, which aligns with our qualitative analysis in the manuscript. Note that, in Table 5, LLE has larger $\Delta y$ than HOPE; however, since LLE has a much larger $y_{0}$ than HOPE, we still identify LLE as the best method within the family of matrix-factorization methods. To make it more visually convenient for the readers, we have visualize Table 5 as Fig. 19. In Fig. 19, points near the bottom-right corner correspond to more robust embedding methods. Again, LLE and node2vec outperform other embedding methods and show best overall robustness.
Figure 19: Visualization of overall robustness metrics ($y_{0}$ and $\Delta y$) for different embedding methods, averaged across all configurations including both random and targeted node selection, various network sizes, mixing parameters, and embedding dimensions.
IV.5 Limitations of Our Study
Having discussed the insights derived from our results, we acknowledge the following limitations of our work:
  1. 1.
    Our experiments focus solely on LFR benchmark graphs, while other benchmarks, such as stochastic block models random graphs, remain unexplored.
  2. 2.
    Our study emphasizes computational investigations of community detection robustness rather than analytical research. Consequently, we do not analyze detectability limits, which remain underexplored for LFR benchmark graphs due to their complex generation and structural properties.
  3. 3.
    In our experiments on two real-world networks, the underlying community structure of ground truth of communities was assigned based on real-world context as determined by domain experts. However, such assignments may introduce subjectivity and potential biases.
  4. 4.
    When we study the two real-world networks, we have artificially reconnecting disconnected components by introducing random edges. This was necessary due to a technical limitation of the graph embedding methods we employed, which tend to perform poorly on disconnected graphs. In fact, the graph embedding methods we used embed each connected component independently for disconnected graphs, leading to arbitrary relative positioning between components, which undermines the global consistency needed for downstream tasks such as clustering or community detection. We acknowledge that artificial reconnection would introduce some deviation from the original network structure.
V Conclusion
@@
Our experimental findings revealed a general decline in community similarity as more nodes were selected and their adjacent edges deleted. Notably, targeted node selection led to a more pronounced and rapid decline in community similarity compared to random node selection. Analysis of clustering similarity scores suggested that LFR networks with lower mixing parameters, indicative of a stronger community structure, exhibit greater robustness to network perturbation. Moreover, diverse graph embedding methods displayed varying degrees of robustness in network community detection. Specifically, LLE demonstrated proficiency within the family of matrix-based methods. For smaller real-world networks with a limited number of vertices, the more robust embedding method is LLE with a modest embedding dimension, typically on the order of tens — such as the 16 dimensions employed in our experiment. In the case of larger real-world networks, the LE and LLE embedding methods showed greater robustness to partition. Specifically, LE with a higher embedding dimension is preferable when the community structure is less pronounced, whereas LE or LLE with a lower embedding dimension is more effective when the network exhibits a stronger community structure. Within the family of random walk methods, our results consistently highlighted node2vec as the superior choice. Notably, node2vec also outperformed all methods within the family of matrix factorization methods.
@@
In contemplating potential avenues for future research, it is noteworthy that our current experiments did not involve parameter optimization in the graph embedding algorithms, primarily due to its inherent computational expense. To attain a more exhaustive and comprehensive insight into the robustness of community detection using graph embedding methods, we may explore the incorporation of parameter optimization in our forthcoming studies. Moreover, while this work focuses on shallow graph embeddings, we acknowledge that neural network-based methods have shown significant promise in recent years. GNNs, particularly in clustering tasks, leverage node features and graph topology in an end-to-end learning framework, providing powerful embeddings that can capture complex relationships within graphs. We plan to explore such approaches in future work to compare their robustness with traditional methods.
@@ Acknowledgements.
Appendix A Remarks on node2vec method
Node2vec produces more robust graph community detection results due to its flexible random walk strategy, which employs biased random walks to provide a trade-off between breadth-first search (BFS) and depth-first search (DFS). As mentioned in [1], node2vec produces higher-quality and more informative embeddings than DeepWalk, enabling node2vec to preserve community structure as well as structural equivalence between nodes.
There has been theoretical evidence supporting the superiority of node2vec over DeepWalk in community detection tasks. As demonstrated in [16, 82], node2vec is notably robust to degree heterogeneity due to its ability to learn degree-agnostic embeddings. In contrast, DeepWalk tends to encode node degree as a dominant feature in the embedding space. This sensitivity makes DeepWalk more susceptible to structural noise in networks with heterogeneous degree distributions, such as those generated by the LFR benchmark model. These theoretical insights are consistent with our empirical results. As shown in Fig. 10 and Fig. 15, when the proportion of selected nodes is small—so the perturbed network retains structural properties close to the original LFR benchmark—the ECS similarity scores achieved by DeepWalk are consistently lower than those by node2vec. A similar pattern is observed in Fig. 11 and Fig. 16. In addition, this theoretical insight is reflected in Tables 4 and 5: the $y_{0}$ values for DeepWalk are smaller than those for node2vec in these two tables.
Furthermore, Cohen et al. [83] studied the robustness of networks with power-law degree distributions under random breakdowns using percolation theory. Specifically, they showed that if a network has a degree distribution with probability mass function $\{P_{k}\}$, and a fraction $p$ of its nodes break down, the resulting degree distribution has probability mass function
$P^{\prime}_{k}=\sum_{\ell=k}^{\infty}P_{\ell}\binom{\ell}{k}(1-p)^{k}p^{\ell-k}.$
[…]
$\operatorname{\mathbf{P}\mathopen{}}\mkern-1.0mu\left[S_{N}>x\right]\sim\operatorname{\mathbf{P}\mathopen{}}\mkern-1.0mu\left[N>\tfrac{x}{1-p}\right],\quad\text{as }x\to\infty.$
Because $N$ is a power-law random variable with exponent $\gamma$ , it follows that
$\operatorname{\mathbf{P}\mathopen{}}\mkern-1.0mu\left[S_{N}>x\right]\sim\bigl{(}\tfrac{x}{1-p}\bigr{)}^{-\gamma},\quad\text{as }x\to\infty.$
Now that $\{P^{\prime}_{k}\}$ is the probability mass function of $S_{N}$, we proved our assertion: if $\{P_{k}\}$ is power-law distributed with exponent $\gamma$, then $\{P^{\prime}_{k}\}$ is asymptotically power-law distributed with the same exponent $\gamma$. Hence, the type of random perturbation we considered do not eliminate degree heterogeneity. Consequently, DeepWalk remains vulnerable to degree-based noise after such perturbations, while node2vec continues to perform robustly in community detection.
As discussed in [16, 67], LINE can be regarded as a special case of node2vec with a window size $w=1$, meaning that it focuses solely on learning dyadic (i.e., direct) relationships between nodes. This limited temporal context restricts LINE’s ability to capture higher-order or indirect structural features. Consequently, LINE exhibits lower robustness than node2vec, particularly in noisy or perturbed networks, because it tends to encode stochastic and noisy pairwise interactions. In contrast, node2vec, by aggregating longer biased random walks, can effectively capture indirect relationships between nodes, making it more resilient in downstream tasks such as community detection.
Moreover, as mentioned in [16], node2vec is robust to network sparsity compared to traditional spectral embedding methods. These observations are corroborated in our experiments. In fact, as we select more nodes and delete their adjacent edges, networks are getting sparsified. Therefore, it is understandable that node2vec is more robust in our experiments.
Hence, node2vec is more robust against random and targeted perturbations in community detection, outperforming other methods in stability under perturbations.
Appendix B M-NMF method with higher embedding dimensions
An increase in ECS can be observed in Figs. 9, 17 and 18 for the M-NMF method, where 16-dimensional and 323232-dimensional embedding are used.
We see some unexpected increasing behaviors in the M-NMF curves in Fig. 9(b) and Fig. 9(d) for synthetic LFR networks with 10,00010,00010 , 000 nodes. Intuitively, these networks are bigger and therefore more complex, which makes it more difficult to detect communities. We thus conducted additional experiments on M-NMF with a 128128128-dimensional embedding as shown in Fig. 20. We observe that the unexpected increasing behaviors are mitigated compared to 323232-dimensional embedding. Specifically, the ECS scores of M-NMF with 128128128-dimensional embedding are more stable than those with 323232-dimensional embedding. This indicates that the choice of embedding dimension can significantly and sensitively impact the performance of M-NMF. Due to the substantial computational cost associated with our additional experiments on synthetic networks with 10,00010,00010 , 000 nodes, we reduced the number of realizations per percentage of selected nodes to 101010, compared to the 505050 used in the original paper, to obtain results within the high-performance computing cluster’s time limit. Standard deviations are still negligible compared to the mean but we do not show them here.
Figure 20: Mean ECS using M-NMF method over the percentage of selected nodes on LFR benchmark graphs with 128128128-dimensional embeddings. (a) and (b) implement random node selection, while (c) and (d) apply targeted selection; (a) and (c) correspond to networks with 1,0001,0001 , 000 nodes, and (b) and (d) to networks with 10,00010,00010 , 000 nodes.
Notably, the M-NMF curves in Fig. 17(a) and Fig. 18(a) for email-EU-core networks are generated using the exact same node selection sequence, with the only difference being the embedding dimension. In Fig. 17(a), where the embedding dimension is 161616, the M-NMF curve remains almost stable showing slightly unexpected increases. In Fig. 18(a), the increasing of ECS is more noticeable. To get more insights into this behavior, we performed additional experiment and analyzed the ECS scores of M-NMF with 128128128-dimensional embedding. The comparison results are presented in Fig. 21. We can see that the ECS scores of M-NMF with 128128128-dimensional embedding are more “stable” as more nodes are selected, where unexpected increasing behaviors are much less pronounced. This again suggests that the choice of embedding dimension can significantly impact the performance of M-NMF, as it has been noticed by [14] and [41]. Furthermore, we observe a difference in community structure: our synthetic LFR networks with 1,0001,0001 , 000 nodes exhibit approximately 101010 communities; our synthetic LFR networks with 10,00010,00010 , 000 nodes exhibit approximately 202020 communities. In contrast, the email-EU-core network contains 424242. Consequently, employing a higher embedding dimension, such as 128128128, appears reasonable to effectively separate these more numerous communities in the embedding space for email-EU-core network.
Figure 21: Mean ECS using M-NMF method on email-EU-core network with 986986986 nodes with different embedding dimensions.
In summary, we have observed that the choice of embedding dimension can significantly sensitively impact the performance of M-NMF in community detection, which can lead to pathological behaviors in the ECS scores. This reveals that M-NMF is not as robust compared to other methods, especially node2vec, which is more stable and less sensitive to the choice of embedding dimension.